OpenStudy (anonymous):

Describe 2 ways in which you can solve -1/2 (5x - 9) = 17.

OpenStudy (anonymous):

-1/2 (5x - 9) = 17 (-2.5x + 4.5) = 17 -2.5x = 12.5 x = -5 -1/2 (5x - 9) = 17 -1 ( -1/2 (5x - 9) = 17) 1/2 (5x - 9) = -17 (2.5x - 4.5) = -17 2.5x = -12.5 x = -5